OpenStudy (anonymous):

Alg 2 Neep help asap Suppose you hit a baseball and its flight takes a pararoblic path. the height of the ball at certain times appears in the "table" below. (Table) Time (s) 0.5 0.75 1 1.25 Height (ft) 10 10.5 9 5.5 A,) Find a quadratic model for the ball's height as a function of time B.) Write athe quadratic function in factored form.

OpenStudy (mathmate):

set up the general quadratic equation (for a parabola) as: y=ax^2+bx+c use 3 of the 4 sets of known points (x,y) to obtain a system of 3 equations, solve for a, b and c. Example: 10=a(0.5)^2+b(0.5)+c 10.5=a(0.75^2+b(0.75)+c 9=a+b+c Solving for a, b and c gives a=-16, b=22, =3 I will leave the factoring to you.

OpenStudy (anonymous):

THANK YOU

OpenStudy (mathmate):

yw! :)